# UniAttn: Reducing Inference Costs via Softmax Unification for Post-Training LLMs

### v1

@@ Abstract
Post-training is essential for adapting Large Language Models (LLMs) to real-world applications. Deploying post-trained models faces significant challenges due to substantial memory overhead and noticeable inference latency. Existing work has identified significant redundancies in LLMs and proposed efficient architectures, namely intra-layer KV sharing and cross-layer KV sharing. However, intra-layer KV sharing still results in high inference costs, while cross-layer KV sharing leads to significant performance degradation. As a result, both methods remain suboptimal for post-training pre-trained LLMs. In this paper, we identify that the Softmax operation is a primary bottleneck for LLM inference and discover that it is actually highly redundant during post-training. We propose Softmax Unification in Attention (UniAttn), a novel post-training method that unifies Softmax activations across transformer blocks to reduce LLM inference costs. Additionally, UniAttn adopts a linear projection to compensate for the errors induced by Softmax unification. Experiments show that UniAttn matches the performance of standard post-training while significantly reducing inference costs, outperforming existing efficient architectures during post-training. Our code will be available at https://github.com/Bostoncake/UniAttn.
@@ 1 Introduction
Figure 1: Comparisons of our UniAttn and directly applying cross-layer KV sharing (CLA) during post-training. “A-X” represents modifying total of X layers when applying method A.
@@
Post-training (Lambert et al., 2024) has become a critical step in developing advanced LLMs for both general-purpose tasks (OpenAI, 2023; Liu et al., 2023) and domain-specific tasks (Thirunavukarasu et al., 2023; Wu et al., 2023; Shao et al., 2024). It typically refers to performing supervised fine-tuning (SFT), reinforcement learning or even domain-specific continual pre-training on pre-trained LLMs using a text corpus different from the original pre-training dataset. Current post-training approaches for real-world applications involve fine-tuning popular base models (Dubey et al., 2024; Jiang et al., 2023; Rivière et al., 2024) with standard decoder-based architectures that have scaled to hundreds of billions of parameters. However, these approaches face two widely recognized challenges in deploying post-trained models (Xiao et al., 2024; Wang et al., 2024a; He et al., 2024; Brandon et al., 2024; Chen et al., 2024): (1) the substantial memory overhead required for KV-cache storage during inference, and (2) the computational latency, which leads to noticeable inference time.
In the literature, researchers have explored methods to achieve efficient inference and identified significant redundancies inside and across LLM layers. This has further led to the development of corresponding efficient architectures: intra-layer Key-Value (KV) sharing, including MQA (Shazeer, 2019) GQA (Ainslie et al., 2023), and cross-layer KV sharing, including CLA (Brandon et al., 2024) and YOCO (Sun et al., 2024). Intra-layer KV sharing methods have been adopted by several open-source base models, including LLaMA-3.1 (Dubey et al., 2024) and Gemma-2 (Rivière et al., 2024), but these models still incur high inference costs. While cross-layer KV sharing has the potential to further reduce inference costs on top of intra-layer KV sharing methods, directly sharing KV-cache across pre-trained LLM layers results in significant performance degradation (Yang et al., 2024). This raises a critical question: Which components in pre-trained LLMs are redundant, and how can they be leveraged to reduce inference costs during post-training while maintaining acceptable performance?
To address this issue, we first conduct a comprehensive analysis of cross-layer KV sharing during post-training. Similar to the training-free scenario (Yang et al., 2024), as shown in Figure 1, cross-layer KV sharing results in substantial performance drops during post-training. To understand this, we then conduct a theoretical analysis and find that cross-layer KV sharing diminishes the impact of model depth on model activations in pre-trained LLMs. Consequently, it undermines the benefits associated with model depth (Wang et al., 2024b), limiting the pre-trained model’s ability to learn new capabilities during post-training. Therefore, alternative approaches beyond cross-layer KV sharing should be explored for post-training.
@@
These findings prompted us to explore other redundancies in pre-trained LLMs, leading to two key observations. First, the Softmax operation significantly contributes to inference costs. Performing the Softmax operation requires access to the entire K-cache. Although the Softmax operation accounts for less than 1% of the FLOPs compared to the linear projections in the backbone, it results in higher latency than the linear projections (Koohpayegani & Pirsiavash, 2024). Second, the Softmax operation does not account for the entire impact of model depth on model activations. Unlike cross-layer KV sharing, we demonstrate in Section 3.4 that eliminating the Softmax operation does not diminish the effect of model depth. Motivated by these observations, we further assess the importance of Softmax operations using a similarity-based metric (Men et al., 2024). Our analysis reveals that Softmax activations exhibit high cross-layer similarity in the top half layers across various open-source pre-trained LLMs and post-training datasets. This observation suggests a generalizable approach to leveraging Softmax redundancies during post-training.
@@
In this paper, we propose to post-train LLMs with Softmax Unification in Attention (UniAttn). Specifically, we group consecutive transformer blocks in the LLM as several “Superblocks”, and unify the Softmax activations of all blocks within each Superblock. This approach significantly reduces memory and inference costs. Additionally, we observe that the error introduced by unifying Softmax activations can be compensated by a linear projection, and in turn design a pipeline to initialize and train the linear projection. We conduct extensive experiments on 4 open-source pre-trained LLMs across two post-training scenarios: enhancing domain-specific capabilities and improving general capabilities. Our results demonstrate that UniAttn consistently achieves performance comparable to standard post-training while reducing inference costs. Compared to directly applying existing efficient architectures to post-training, UniAttn achieves substantially better performance with lower inference costs. Furthermore, our UniAttn can be combined with KV-Cache compression methods to further cut down the memory overhead, showing strong practicality for real-world applications.
@@
  • We consider a critical and practical challenge of leveraging the redundancies in pre-trained LLMs for post-training. We show theoretically and experimentally that existing methods are suboptimal.
  • We study the redundancies in pre-trained LLMs and discovered significant redundancies in the Softmax operation. We propose Softmax Unification in Attention (UniAttn) for post-training LLMs. Specifically, we unify the Softmax activations across blocks inside the LLM. Additionally, we leverage linear projection to compensate for the error from Softmax unification.
  • Extensive experiments show that UniAttn achieves comparable performance to standard post-training while reducing inference costs, and outperforms existing efficient architectures as well.
@@ 2.1 Post-Training LLMs
Building frontier LLMs for real-world applications involves two crucial stages: pre-training and post-training. Since pre-training data and methodologies are often proprietary, the research community has extensively explored post-training upon open-source pre-trained LLMs (Dubey et al., 2024; Rivière et al., 2024; Jiang et al., 2023) to enhance their general or domain-specific capabilities for deployment (Lian et al., 2024). Post-training is typically performed on instruction-following or domain-specific datasets. Recently, various datasets have been curated to equip LLMs with specific abilities, including general instruction-following models (Teknium et al., 2024; Lambert et al., 2024; BAAI, 2024), medical QA models (Wu et al., 2023; Gururajan et al., 2024; Singhal et al., 2023), legal QA models (Huang et al., 2023; Zhou et al., 2024), and models with strong mathematical problem-solving capabilities (Liu & Low, 2023).
@@
2.2 Efficient LLM Architectures
Efficient architectures utilize structural redundancies to create inference efficient model variants (Bolya et al., 2023; Xiong et al., 2024) for deployment. In the scope of LLMs, efficient architectures mainly fall into two categories: intra-layer KV sharing, including MQA (Shazeer, 2019) and GQA (Ainslie et al., 2023), and cross-layer KV sharing, including CLA (Brandon et al., 2024) and YOCO (Sun et al., 2024). Specifically, MQA (Shazeer, 2019) simplifies the attention mechanism by utilizing multiple query heads and a single KV head. GQA (Ainslie et al., 2023) takes a step further from MQA by organizing query heads as multiple groups and assigns different KV heads to each group. CLA (Brandon et al., 2024) proposes a cross-layer sharing mechanism to further reduce KV-cache memory overhead by sharing KV-cache across different layers. YOCO (Sun et al., 2024) transforms the original structure into self-decoders and cross-decoders, and adopts a global KV-cache across decoders.
Existing intra-layer KV sharing works have been adopted by various open-source LLMs (Dubey et al., 2024; Jiang et al., 2023; Rivière et al., 2024). However, due to their inherent massive scale, those models still incur significant inference costs. According to our analysis in Figure 1 and Section 3.2, directly applying cross-layer KV sharing for post-training is suboptimal. To address this, we propose UniAttn, which achieves promising performance for post-training LLMs and outperforms competing methods.
@@ 3.1 Preliminaries
In this paper, we focus on mainstream decoder-only LLMs with a transformer structure. Each transformer block in an LLM features a multi-head self-attention (MHA) module and a feed-forward network (FFN). Given an input sequence $\mathbf{x}\in\mathbb{R}^{l\times d}$ where $l$ denotes the sequence length and $d$ denotes the hidden size, both MHA and FFN generate an output sequence with identical length and shape. We focus on the MHA module. Formally, we denote MHA output in layer $i$ as $\mathbf{x}_{i}^{\prime}$, where $\mathbf{x}_{i}^{\prime}=\text{MHA}(\text{Norm}(\mathbf{x}_{i}))+\mathbf{x}_{i}$. “$\text{Norm}(\cdot)$” denotes the Pre-Norm, a component adopted by mainstream open-source LLMs (Dubey et al., 2024; Jiang et al., 2023; Rivière et al., 2024; Bai et al., 2023). In the MHA module, each token in the input sequence $\mathbf{x}$ is first being projected by $W_{Q}$, $W_{K}$, and $W_{V}$, forming $Q,K\in d\times d_{k}$ and $V\in d\times d_{v}$. Then, the Softmax activation $s_{i}$ is calculated by:
$s=\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}}).$ (1)
@@
Subsequently, $s$ is projected using $V$ and the weight matrix $W_{o}$, and the input to MHA is added back through a residual connection to produce the MHA output $\mathbf{x}_{i}^{\prime}$:
$\mathbf{x}_{i}^{\prime}=sVW_{o}+\mathbf{x}_{i}$ (2)
3.2 Analysis on Existing Methods
To address redundancies in LLMs, mainstream efficient LLM architectures typically remove insignificant modules in the backbone. Among these methods, intra-layer KV sharing methods such as GQA (Ainslie et al., 2023) has been widely adopted by numerous open-source pre-trained models, like LLaMA3, Mistral, and Gemma2. On the other hand, cross-layer KV sharing methods, including CLA (Brandon et al., 2024) and YOCO (Sun et al., 2024), share the same KV-cache across different layers. Through theoretical analysis, we show that cross-layer KV sharing diminishes the impact of model depth on model activations. This negatively affects the capabilities of pre-trained models, as model depth is a critical factor in determining their overall performance (Wang et al., 2024b).
We first give a series of conclusions as preliminaries.
Lemma 3.1.
$f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, $i\in\mathbb{N}$ are LINEAR transformations and $\text{Norm}(\cdot)$ yields a vector with unit Frobenius-norm (denoted as |||||$\cdot$||| | ⋅ | |). Let $\mathbf{x}_{0}\in\mathbb{R}^{n}$ be the input to the system. Consider an $L$-layer Pre-Norm architecture defined by:
$\mathbf{x}_{k}=\mathbf{x}_{k-1}+f_{k}(\text{Norm}(\mathbf{x}_{k-1})),\quad\text{for}\quad k=1,2,\dots,L.$ (3)
$||\mathbf{x}_{L}||\leq||\mathbf{x}_{0}||+\lambda L$ (4)
See Appendix for the proof. Lemma 3.1 shows that a Pre-Norm linear system has bounded growth. The norm of output by each layer grows at most linearly with depth $L$.
Proposition 3.2.
Pre-trained decoder-based LLMs exhibit a high linearity score. Formally, let $A,B\in\mathbb{R}^{n\times d}$ denote the normalized input and output of a decoder block in LLM, respectively,
$\min_{X}||AX-B||\approx 1$ (5)
Proposition 3.2 has been validated by (Razzhigaev et al., 2024) on pre-trained LLMs, suggesting that we can approximate each decoder block by a linear transformation with bounded singular values, thus placing pre-trained LLMs structurally “close” to a linear Pre-Norm system as in Lemma 3.1. In the following discussion, we approximate pre-trained LLMs with an equivalent linear Pre-Norm system and investigate its input and output in each layer.
Proposition 3.3.
The input and output of a layer in pre-trained decoder-based LLMs share a high cosine similarity.
Proposition 3.3 has been validated by (Men et al., 2024) on pre-trained LLMs, indicating that each layer in LLMs predominantly change the magnitude of the activation. We leverage the following conclusion for analyzing the change of activation magnitude.
Proposition 3.4.
Training Pre-Norm transformers leads to top layers receiving smaller updates (gradient norms decrease from bottom to top).
(Xiong et al., 2020) has given both a theoretical proof sketch and experimental evidence of Proposition 3.4. Together with Proposition 3.2, we can reasonably assume that the largest singular values of the top layers tend to be smaller after training (See Appendix for demonstration). Since pre-trained LLMs mainly change the magnitude of the activation (as shown in Proposition 3.3), we can use the size of the largest singular values in each layer as an indicator of the corresponding activation magnitude change. This leads to the assumption that the top-layer transformations generally generate outputs with smaller norms.

### v2

@@ Abstract
Post-training is essential for adapting Large Language Models (LLMs) to real-world applications. Deploying post-trained models faces significant challenges due to substantial memory overhead and noticeable inference latency. Existing work has identified significant redundancies in LLMs and proposed efficient architectures, namely intra-layer KV sharing and cross-layer KV sharing. However, these methods still result in high inference time overhead, remaining suboptimal for post-training pre-trained LLMs. In this paper, we identify that the Softmax operation is a primary bottleneck for LLM inference and discover that it is actually highly redundant during post-training. We propose Softmax Unification in Attention (UniAttn), a novel post-training method that unifies Softmax activations across transformer blocks to reduce LLM inference costs. Additionally, UniAttn adopts a linear projection to compensate for the errors induced by Softmax unification. Experiments show that UniAttn matches the performance of standard post-training while significantly reducing inference costs, outperforming existing efficient architectures during post-training.
UniAttn, Large Language Models, Softmax
@@ 1 Introduction
Post-training (Lambert et al., 2024) has become a critical step in developing advanced LLMs for both general-purpose tasks (OpenAI, 2023; Liu et al., 2023) and domain-specific tasks (Thirunavukarasu et al., 2023; Wu et al., 2023; Shao et al., 2024). It typically refers to performing supervised fine-tuning (SFT) or domain-specific continual pre-training on pre-trained LLMs using a text corpus different from the original pre-training dataset. Current post-training approaches for real-world applications involve fine-tuning popular base models (Dubey et al., 2024; Jiang et al., 2023; Rivière et al., 2024) with standard decoder-based architectures that have scaled to hundreds of billions of parameters. However, these approaches face two widely recognized challenges in deploying post-trained models (Xiao et al., 2024; Wang et al., 2024a; He et al., 2024; Brandon et al., 2024; Chen et al., 2024): (1) the substantial memory overhead required for KV-cache storage during inference, and (2) the computational latency, which leads to noticeable inference time.
To address those challenges, recent research has proposed KV-cache sharing methods, and showed that they offer notable advantages over traditional pruning (Ma et al., 2023; Zhang et al., 2023) and quantization (Xiao et al., 2023; Lin et al., 2024) methods: While pruning maintains model accuracy with only modest efficiency gains, and quantization reduces inference overhead at the expense of substantial performance degradation, KV-cache sharing methods achieve efficiency improvements with minimal performance drops. To this end, researchers have identified significant redundancies inside and across LLM layers, leading to two mainstream KV-cache sharing paradigms: intra-layer KV sharing, including MQA (Shazeer, 2019) and GQA (Ainslie et al., 2023), and a more aggressive cross-layer KV sharing, including CLA (Brandon et al., 2024) and YOCO (Sun et al., 2024). Despite achieving impressive KV-cache compression ratios, cross-layer KV sharing yields limited inference time reduction: less than 5% at deployment (see Section 4 for detailed results). This raises a critical question: How to design LLM architectures that achieve both inference time and memory efficiency?
@@
To further explore that question, we delve deeper into LLM architectures and make two key observations. First, the Softmax operation significantly contributes to inference costs. Performing the Softmax operation requires access to the entire K-cache. Although the Softmax operation accounts for less than 1% of the total FLOPs compared to the linear projections in the backbone, it results in higher latency than the linear projections due to limited parallelism (Koohpayegani and Pirsiavash, 2024; Liu et al., 2024). Second, Softmax activations exhibit high cross-layer similarity in the top half layers across various open-source pre-trained LLMs and post-training datasets. For instance, the average cross-layer similarity exceeds 0.9 on the top half layers in LLaMA-3.1 8B (Dubey et al., 2024) when evaluated on the Tulu3 (Lambert et al., 2024) training set (see Figure 2 for detailed results). This observation highlights the potential for an effective and generalizable approach to leveraging Softmax redundancies during post-training.
@@
In this paper, we propose to post-train LLMs with Softmax Unification in Attention (UniAttn). Specifically, we group consecutive transformer blocks in the LLM as several “Superblocks”, and unify the Softmax activations of all blocks within each Superblock. This approach significantly reduces memory and time costs. Additionally, we demonstrate that the error introduced by unifying Softmax activations can be effectively compensated by a linear projection, for which we design an initialization and training pipeline. To better interpret the performance gains of UniAttn, we present a theoretical analysis comparing it with existing cross-layer KV sharing methods. Our analysis reveals that UniAttn better preserves the benefits associated with model depth (Wang et al., 2024b), enabling the model to acquire new capabilities during post-training.
Figure 1: Comparisons of our UniAttn and directly applying cross-layer KV sharing (CLA) during post-training. “A-X” represents modifying total of X layers when applying A.
We conduct extensive experiments on four open-source pre-trained LLMs across two post-training scenarios: enhancing domain-specific capabilities and improving general capabilities. Our results show that UniAttn consistently achieves performance comparable to standard post-training while reducing inference costs. Compared to existing KV-cache sharing baselines, UniAttn achieves better performance with substantially lower inference costs (see Figure 1 for an overview). Furthermore, our UniAttn can be combined with intra-layer KV sharing and KV-cache compression methods to further cut down the memory overhead, showing strong practicality for real-world applications.
@@
  • We identify that the Softmax operations in pre-trained LLMs, although yielding high inference costs, exhibit significant redundancies as Softmax activations share high cross-layer similarities.
  • Based on the observations, we propose Softmax Unification in Attention (UniAttn) for post-training LLMs. Specifically, we unify the Softmax activations across decoder blocks. Additionally, we leverage linear projection to compensate for the error from Softmax unification. We show theoretically that UniAttn better preserves model capabilities than compared baselines.
  • Extensive experiments show that UniAttn achieves comparable performance to standard post-training while reducing inference costs, outperforming existing efficient architectures.
@@ 2.1 Post-Training LLMs
Building frontier LLMs for real-world applications involves two crucial stages: pre-training and post-training. Since pre-training data and methodologies are often proprietary, the research community has extensively explored post-training upon open-source pre-trained LLMs (Dubey et al., 2024; Rivière et al., 2024; Jiang et al., 2023) to enhance their general or domain-specific capabilities for deployment (Lian et al., 2024). Post-training is typically performed on instruction-following or domain-specific datasets. Recently, various datasets have been curated to equip LLMs with specific abilities, including general instruction-following models (Teknium et al., 2024; Lambert et al., 2024; BAAI, 2024), medical QA models (Wu et al., 2023; Gururajan et al., 2024; Singhal et al., 2023), legal QA models (Huang et al., 2023; Zhou et al., 2024), and models with strong mathematical problem-solving capabilities (Liu and Low, 2023).
@@
2.2 Efficient KV Sharing Architectures
Unlike existing pruning and quantization approaches (Bolya et al., 2023; Xiong et al., 2024), efficient KV sharing architectures utilize structural redundancies to create inference-efficient model variants for deployment. Those architectures mainly fall into two categories: intra-layer KV sharing, including MQA (Shazeer, 2019) and GQA (Ainslie et al., 2023), and cross-layer KV sharing, including CLA (Brandon et al., 2024) and YOCO (Sun et al., 2024). Specifically, MQA (Shazeer, 2019) simplifies the attention mechanism by utilizing multiple query heads and a single KV head. GQA (Ainslie et al., 2023) takes a step further from MQA by organizing query heads as multiple groups and assigns different KV heads to each group. CLA (Brandon et al., 2024) proposes a cross-layer sharing mechanism to further reduce KV-cache memory overhead by sharing KV-cache across different layers. YOCO (Sun et al., 2024) transforms the original structure into self-decoders and cross-decoders, and adopts a global KV-cache across decoders.
Existing intra-layer KV sharing works have been adopted by various open-source LLMs (Dubey et al., 2024; Jiang et al., 2023; Rivière et al., 2024), yielding modest reduction on inference memory overhead. Although cross-layer KV sharing introduces more aggressive memory reduction, according to our analysis in Section 1 and Section 3.3, it still suffers from efficiency and performance issues. To address this, we propose UniAttn, which achieves promising performance for post-training LLMs and outperforms competing methods.
@@ 3.1 Preliminaries
In this paper, we focus on mainstream decoder-only LLMs with a transformer structure. Each transformer block in an LLM features a multi-head self-attention (MHA) module and a feed-forward network (FFN). Given an input sequence $\mathbf{x}\in\mathbb{R}^{l\times d}$ where $l$ denotes the sequence length and $d$ denotes the hidden size, both MHA and FFN generate an output sequence with identical length and shape. We focus on the MHA module. Formally, we denote MHA output in layer $i$ as $\mathbf{x}_{i}^{\prime}$, where $\mathbf{x}_{i}^{\prime}=\text{MHA}(\text{Norm}(\mathbf{x}_{i}))+\mathbf{x}_{i}$. $\text{Norm}(\cdot)$” denotes the Pre-Norm, a component adopted by mainstream open-source LLMs (Dubey et al., 2024; Jiang et al., 2023; Rivière et al., 2024; Bai et al., 2023). In the MHA module in layer $i$, each token in the input sequence $\mathbf{x}_{i}$ is first projected by $W_{Q,i}$, $W_{K,i}$, and $W_{V,i}$, forming $Q_{i},K_{i}\in d\times d_{k}$ and $V_{i}\in d\times d_{v}$. Then, the Softmax activation $s_{i}$ is calculated by:
$s_{i}=\text{softmax}(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{k}}}).$ (1)
@@
Subsequently, $s_{i}$ is projected using $V_{i}$ and the output weight matrix $W_{o,i}$ of the $i$th layer, and the input to MHA is added back through a residual connection to produce the MHA output $\mathbf{x}_{i}^{\prime}$:
$\mathbf{x}_{i}^{\prime}=s_{i}V_{i}W_{o,i}+\mathbf{x}_{i}$ (2)
3.2 UniAttn: Softmax Unification in Attention
Although existing KV-cache sharing methods–such as intra-layer and cross-layer KV sharing–can reduce the memory footprint, they offer limited inference time reduction. To achieve both memory-efficiency and inference acceleration, we investigate the Softmax operation because (1) calculating Softmax demands the entire K-cache that accounts for 50% memory of the KV-cache; and (2) multiple studies (Geng et al., 2018; Koohpayegani and Pirsiavash, 2024) have recognized that the Softmax operation leads to high inference latency. We study the redundancy of Softmax activations in pre-trained LLMs by measuring the importance of Softmax operations in each layer. To ensure the generalizability of our study, we employ post-training datasets PMC (Wu et al., 2023), Infinity Instruct (Math) (BAAI, 2024), Tulu3 (Lambert et al., 2024), and long samples (F15K) provided by (Su, 2023), and utilize LLaMA-2 7B (Touvron et al., 2023), LLaMA-3.1 8B (Dubey et al., 2024), Mistral 7B (Jiang et al., 2023), and Gemma-2 9B (Rivière et al., 2024). We collect all Softmax activations, i.e., the $s_{i}$ vectors in Equation 1, and compute the cosine similarity $\text{Sim}(i)=\cos(s_{i},s_{i-1})$ for each sequence. We then average $\text{Sim}(i)$ over all sequences.
Observation: As shown in Figure 2, Softmax activations in top half layers share a high cross-layer similarity across different settings, showing that Softmax operations are highly redundant in LLMs. See Appendix C for quantitative analysis.
Figure 2: Cosine similarity results of average Softmax activations. Across all settings, the average Softmax activations of the top half of layers (i.e., the right half of each heatmap bar) share a high cosine similarity.
Motivated by such redundancy, we propose Softmax Unification in Attention (UniAttn) for post-training LLMs. Specifically, UniAttn groups consecutive decoder blocks in LLMs as “Superblocks”, and unifies the Softmax activations inside each Superblock to reduce the full Softmax operations and KV-cache size during inference. Additionally, we observe that the error introduced by Softmax unification can be compensated by a linear projection, and in turn design a pipeline to initialize and train the linear projection. UniAttn simultaneously reduces the GPU memory and inference latency while achieving competitive performance. Different from cross-layer KV sharing, UniAttn does not diminish the “depth factor”. We show this in Section 3.4.
$\mathbf{x}_{i+b}^{\prime}=s_{i}V_{i+b}W_{o,i+b}+\mathbf{x}_{i+b}$ (3)
Since only the bottom block contribute to calculating the Key projection, other blocks that reuse Softmax activations do not store their corresponding K-cache, leading to GPU memory-efficiency. Moreover, removing Softmax calculations in those layers also contribute to inference latency reduction. Similar to cross-layer KV sharing, UniAttn is also orthogonal to intra-layer KV sharing. In Section 4, we employ pre-trained LLMs with GQA to evaluate UniAttn.
Linear Compensation. Simply unifying Softmax activations unavoidably introduces errors in the forward pass. Formally, the error $\epsilon$ in layer $i+b$ is:
$\epsilon$ $=\mathbf{x}_{i+b}^{\text{ori}\ \prime}-\mathbf{x}_{i+b}^{\text{uni}\ \prime}=(\mathbf{x}_{i+b}^{\text{ori}}-\mathbf{x}_{i+b}^{\text{uni}})$ (4)
$+(\text{softmax}(\frac{Q_{i+b}K^{T}_{i+b}}{\sqrt{d_{k}}})-s_{i})V_{i+b}W_{o,i+b}$
where $\mathbf{x}_{i+b}^{\text{ori}}$ and $\mathbf{x}_{i+b}^{\text{Uni}}$ denote the input to layer $i+b$ in the original model and the model with UniAttn applied, respectively. According to (Koohpayegani and Pirsiavash, 2024), calculating linear projections yield significantly lower latency compared to the Softmax operation. And thus we study the effectiveness of compensating $\epsilon$ with a linear projection matrix $W_{c}$. To compensate for $\epsilon$, we apply $W_{c}$ to each layer with unified Softmax activations as:
$\mathbf{x}_{i+b}^{\prime}=s_{i}V_{i+b}W_{o,i+b}+\mathbf{x}_{i+b}+\mathbf{x}_{i+b}W_{c,i+b}$ (5)
To find a proper initialization for $W_{c}$ in a training-free manner, we propose to directly compensate $\epsilon$:
Theorem 3.1.
The initialization for $W_{c}$ that incurs minimal error when compensating $\mathbb{E}(\epsilon)$ satisfies:
$W_{c}=V\Sigma^{+}U^{T}\mathbb{E}(\epsilon),$ (6)
We leave the proof in the Appendix. In practice, we forward a small portion of training data from the post-training datasets, and use the average values $\bar{\mathbf{x}}_{i+b}$ and $\bar{\epsilon}$ as estimations of $\mathbb{E}(\mathbf{x}_{i+b})$ and $\mathbb{E}(\epsilon)$. Considering that compensating for errors in previous layers would influence the calculation in subsequent layers, we calculate the initialization of $W_{c}$ bottom-up. Experiments on the F15K dataset (Su, 2023) with sequence length 5,120 show that, after linear compensation, the errors for LLaMA-2 7B and LLaMA-3.1 8B are reduced to 2.9% (11.09/382.87) and 6.4% (5.76/89.58) of the original errors without compensation, respectively. This clearly demonstrates the effectiveness of the linear compensation. Please refer to Appendix I for more theoretical insights on the linear compensation.
To better train the linear transformations, we adopt a two-stage post-train pipeline. In the first stage, we only fine-tune the $W_{c}$ weights with early stop and keep other weights frozen. We then conduct a full fine-tuning in the second stage. We formalize a detailed pipeline in Algorithm 1 in the Appendix.
Figure 3: Pipeline comparison between standard decoder-based transformers, CLA (Brandon et al., 2024) (block size 3), and UniAttn (Superblock size 3). UniAttn shares the Softmax activations across layers in grouped Superblocks and adds linear transformation $W_{c}$ to compensate for the unification error. For simplicity, only the self-attention calculation is illustrated for CLA and UniAttn.
3.3 Theoretical Analysis on Existing Methods
Existing KV-cache sharing methods can be mainly categorized as intra-layer KV sharing and cross-layer KV sharing methods. Intra-layer KV sharing methods can be directly combined with UniAttn (see Section 4 for detailed results). While cross-layer KV sharing methods achieve a more aggressive memory reduction rate, through theoretical analysis, we show that it diminishes the impact of model depth on model activations through directly sharing the same KV-cache across layers. This negatively affects the capabilities of pre-trained models, as model depth is a critical factor in determining their overall performance (Wang et al., 2024b).
Existing research (Razzhigaev et al., 2024) has shown that LLM layers can be approximated as linear transformations. Together with the conslusion on training gradient dynamics (Xiong et al., 2020), we first give a reasonable assumption: